OpenStudy (vshiroky):

Find an equation of the line has the given slope and passes through the given point. m=-3 (1/2, 3)

OpenStudy (mertsj):

Why don't you use point-slope form?

OpenStudy (vshiroky):

I think it's y=-3x+3

OpenStudy (mertsj):

\[y-3=-3(x-\frac{1}{2})\]

OpenStudy (mertsj):

Simplify that.

OpenStudy (vshiroky):

I tried that's what I came up with

OpenStudy (mertsj):

Show me your work. I will find your mistake.